### v1

# A Comparative Analysis of Wealth Index Predictions in Africa between three Multi-Source Inference Modelsthanks: Please cite the ECMLPKDD’24 version of this paper

Abstract
Poverty map inference is a critical area of research, with growing interest in both traditional and modern techniques, ranging from regression models to convolutional neural networks applied to tabular data, images, and networks. Despite extensive focus on the validation of training phases, the scrutiny of final predictions remains limited. Here, we compare the Relative Wealth Index (RWI) inferred by Chi et al. (2021) with the International Wealth Index (IWI) inferred by Lee and Braithwaite (2022) and Espín-Noboa et al. (2023) across six Sub-Saharan African countries. Our analysis focuses on identifying trends and discrepancies in wealth predictions over time. Our results show that the predictions by Chi et al. and Espín-Noboa et al. align with general GDP trends, with differences expected due to the distinct time-frames of the training sets. However, predictions by Lee and Braithwaite diverge significantly, indicating potential issues with the validity of the model. These discrepancies highlight the need for policymakers and stakeholders in Africa to rigorously audit models that predict wealth, especially those used for decision-making on the ground. These and other techniques require continuous verification and refinement to enhance their reliability and ensure that poverty alleviation strategies are well-founded.
1 Introduction
Accurate and detailed poverty maps are indispensable for targeting anti-poverty interventions. While the United Nations prioritizes poverty eradication [25], official poverty maps often do not exist [21] or lack the granularity to identify specific areas of critical need due to data aggregation at broad administrative levels [24]. In recent years, predicting poverty using novel data sources has garnered significant attention within the AI for social good community. Diverse approaches leveraging satellite imagery [10, 13], mobile phone data [4, 6], social media activity [17, 19], or combinations thereof [2, 16, 8], have emerged, raising questions about the optimal methodology for accurate poverty estimation. Towards this goal, Espín-Noboa et al. 2023 [8] examined the performance of three machine learning architectures utilizing satellite imagery, tabular data, and a hybrid approach. Their findings revealed a nuanced landscape, with image-based models excelling at inferring the wealthiest areas, while tabular features proved more effective in rural regions. These findings highlight the importance of tailoring models to specific regions and feature sources.
However, current methods for creating poverty maps have some limitations. Although model architectures, features, and train-test validations are meticulously designed and evaluated on small samples, they may not ensure accurate predictions in the final poverty maps. The lack of ground-truth data prevents the validation of these predictions, leading to potential inaccuracies. Consequently, these unvalidated predictions are used for policy-making, relying on training samples that may not accurately represent the entire population.
Another critical issue is the choice of ground-truth indicators. Various studies use socioeconomic indicators such as income per capita [7], housing prices [26], and wealth indexes [2, 16, 8], each with its strengths and limitations [20]. Different wealth indexes such as the Relative Wealth Index (RWI), the Comparative Wealth Index (CWI), the Harmonized Wealth Index (HWI), and the International Wealth Index (IWI), provide distinct perspectives on wealth but also come with inherent constraints [16]. This variability makes it challenging to determine the most reliable inferred poverty map when comparing methodologies with similar predictive power.
If the validations of these different models are equally accurate, their resulting poverty maps should theoretically be similar. However, any discrepancies would prompt an investigation into the reasons for these differences and an assessment of which model is the most accurate. In this study, we aim to address this gap by comparing three recent wealth inference approaches: RWI by Chi et al. [2], and IWI by Lee and Braithwaite [16], and by Espín-Noboa et al. [8]. We show their inferred poverty maps across six African countries and examine the degree of concordance of these predictions with established GDP trends, aiming to evaluate the robustness and reliability of these methodologies. The analysis is two-fold. First, we compare the predicted distributions of wealth to assess the overall poverty inferred for each country by each model. Second, considering the differences in geo-located places each approach uses in the final poverty map, we verify whether overlapping locations across methods receive consistent predictions and align with expected poverty trends. In adherence to open science practices, we make both the code and the predictions openly available [9].
2 Preliminaries
In this study, we compare the predictions of high-resolution poverty maps across six African countries from three different studies: M1 by Chi et al. (2021) [2], M2 by Lee and Braithwaite (2022) [16], and M3 by Espín-Noboa et al. (2023) [8].
2.1 Ground-Truth Data
Ground-truth data in M1, M2, and M3 are derived from nationally-representative, population-based household surveys conducted by the Demographic and Health Survey (DHS) Program [12]. These surveys encompass asset/wealth information and sub-regional geomarkers. Each household survey includes numerous questions on socioeconomic status. From the responses, DHS calculates the Relative Wealth Index (RWI) [3, 23] using the first principal component derived from a Principal Component Analysis (PCA) [11, 14] of a standardized set of 15 questions concerning assets and housing characteristics. These questions cover various aspects such as electricity, telephone, automobile, refrigerator, TV, water supply, toilet facilities, and construction materials. This RWI serves as the ground-truth measure of wealth in M1. In contrast, M2 and M3 compute the International Wealth Index (IWI) [22], a similar metric, which assesses household material well-being and economic status across low and middle-income countries. The IWI is derived from 12 housing characteristics; M2 excludes two of them, while M3 uses them all. PCA is used to determine asset weights, which are re-scaled to produce IWI scores ranging from 0 to 100, representing the spectrum from no assets to the highest wealth. The IWI’s advantage lies in its comparability across different countries and years, unlike the country-specific RWI used in M1.
The studies also differ in their spatial resolution. M1 employs a 2.4 km grid system defined by Bing tile coordinates to map DHS clusters to grid cells. Conversely, M2 and M3 use actual DHS cluster locations which typically contain 25-30 households, and are geo-located with added noise to protect respondents privacy. The noise varies: up to 2 km for urban and 5 km for rural clusters, with 1% of rural clusters displaced by up to 10 km [1].
Table 1: Years of collected DHS ground-truth and features per study. Instances where two or more models used the same ground-truth data are few: M2 and M3 in Uganda, M1 and M2 in Rwanda, and all models in South Africa. M2 inferred Gabon’s poverty map using a model trained on data from 25 Sub-Saharan African countries, as the most recent survey from 2012 was considered outdated. The table also categorizes the features utilized by each model. Note that for this analysis, as compared to the original models reported in [8], the models of M3 were retrained without features from mobility and demographics, as the mobility data was discontinued and the Facebook Marketing API redefined its reach estimates. This change did not affect performance.
Country M1 M2 M3
Sierra Leone 2013 2016 2016,2019
Liberia 2013 2016 2019,2022
Uganda 2016 2016,2018-19 2016,2018-19
Rwanda 2014-15 2014-15 2019
South Africa 2016 2016 2016
Gabon 2012 Multiple 2019
Feature M1 M2 M3*
Daylight satellite images
Nightlight intensity
Infrastructure
Connectivity -
Population
Mobility network $\dagger$ - - (✓)
Demographics of people $\dagger$ - - (✓)
Natural conditions - -
Number of total features $\ddagger$ 12+100 249+4 173+784
* This model has three variants: CNN (only images), CatBoost (all but images), and combined (all features).
$\dagger$ We exclude these features due to discontinued support (mobility) and rapid changes of API (demographics).
$\ddagger$ Metadata (non-imagery) features (12, 249, and 173, resp.) plus image-based features (100, 4, and 784, resp.).
The final poverty map is defined by the grid-cells covering the entire country (excluding those used in training) in M1, and by the populated places (i.e., cities, towns, villages, hamlets and isolated dwellings) extracted from OpenStreetMap (OSM) in M2 and M3 (M3 additionally includes neighborhoods).
Another critical difference among the three studies is the year in which the surveys were conducted. Each study utilized training data from different years (see Table 1), complicating the direct comparison of their results. Consequently, as explained in Section 2.3, our comparison will not concentrate on the precise predictions of each study. Instead, we will focus on identifying expected trends, using the Gini coefficient of wealth distributions and GDP per capita as baselines. These baselines are sourced from official data provided by the World Bank [27, 28]. This approach allows us to account for temporal variations and provides a more meaningful analysis of the wealth and economic status trends across different periods.
Figure 1: Available GPD per capita and Gini index. Top: GDP values are aligned with the years of the surveys used in the models (x-axis). Bottom: The most recent available Gini indices are shown for reference. Uganda is the only country with GDP and Gini index data available for approximately the same years as the surveys. Both metric values are provided by the World Bank.
2.2 Models
M1: RWI inference model by Chi et al. 2021 [2, 5] utilizes $112$ features in total, including high-resolution satellite imagery, mobile phone network data, topographic maps, and aggregated, de-identified connectivity data from Facebook. These diverse data sources are processed through deep learning and computational algorithms to generate quantitative features for each village. These features are then used to train a gradient-boosted regression tree, a supervised machine-learning model, to predict RWI scores. Hyperparameters for the gradient-boosted tree are tuned to minimize the cross-validated mean-squared error (MSE), using a grid search. They use three cross-validation approaches: 5-fold cross-validation (basic CV), leave-one-country-out cross-validation (leave-country-out), and spatially stratified cross-validation (spatial CV). The use of different cross-validation methods helped highlight the potential upward bias in performance due to spatial auto-correlation in the training and test data. All analysis in this study is based on spatial CV, ensuring conservative and appropriate model training for geographic data with spatial auto-correlation. For our comparison analysis, we linearly transform the RWI scores to their IWI equivalents in M3 (see Equation 1) as both are highly correlated (i.e., Pearson’s correlation coefficients: $0.93-0.99$ [16]).
$\hat{iwi}=\frac{(rwi-rwi_{min})*(iwi_{max}-iwi_{min})}{(rwi_{max}-rwi_{min})}+iwi_{min}$ (1)
M2: IWI inference model by Lee and Braithwaite 2022 [16] combines feature-based and image-based models to enhance each other’s performance. It uses two cross-validation approaches similar to M1: 5-fold cross-validation (basic CV) and leave-one-country-out cross-validation (leave-country-out). Initially, an XGBoost (eXtreme Gradient Boosting) algorithm is trained on $249$ features (infrastructure, night-time luminosity, and population density) to predict IWI scores. Bayesian optimization is employed to tune the hyperparameters, selecting the best configuration for the training dataset. The XGBoost model’s predictions are then used to train a Convolutional Neural Network (CNN) model that outputs a probability distribution of socioeconomic status (rich, upper-middle class, lower-middle class, poor) for a 1-square-mile area. These probabilities are then fed back into the XGBoost model, creating a cycle where the models iteratively improve by providing better training data.
M3: IWI inference model by Espín-Noboa et al. 2023 [8] extracts $957$ features from seven data sources. Each location is defined by the centroid of each place, using a default bounding-box width of 1 mile. Some features are also queried within 2, 5, and 10 Km to capture the original locations of the DHS clusters, similar to M2. Features include high-resolution satellite imagery, demographics, mobility networks, and mobile phone antennas. Three machine learning models are proposed to predict both the mean and standard deviation of IWI scores, providing insights into the wealth distribution within clusters. The first model is a CNN that predicts IWI scores from daylight satellite images. This CNN comprises 22 layers, with the final layer using a linear activation function and MSE as the loss function, tuning 4 hyperparameters. The second model is a CatBoost regressor (CB), predicting IWI scores from 173 metadata features and evaluated individually for each data source. This model is tuned using 11 hyperparameters. The third model combines CNN+CB by feeding the third-to-last layer of the CNN into the CB model, adding 784 features. For training, data is partitioned into 80% train and 20% test sets, stratified into 10 equal-width bins of wealth. A 4-fold cross-validation on the train set, stratified by SES (ses CV), is used to tune hyperparameters via Random Search on 200 combinations. This process is repeated three times independently with different random seeds, to control for fluctuations, and mean performance is reported. The models are evaluated on various configurations of training data, including different years of ground-truth data, relocation strategies, data augmentation ($*_{a}$), sample weight strategies ($*_{w}$), and feature sources, ensuring robust and comprehensive performance evaluation. For this analysis, we choose the model with best performance (lowest root-mean-square error RMSE) in each country: CB for Sierra Leone and Liberia, CBw for Uganda and Rwanda, and CNNa+CB for Gabon and South Africa.
In summary, all models use a multi-modal and hybrid approach, combining features from various data sources and two model architectures, see Table 1 (bottom). M1 and M3 use satellite images by pre-training a CNN model and feeding the image embeddings into a gradient boosting algorithm. M2 also uses a hybrid approach but first pre-trains a gradient boosting algorithm to train a CNN, and then creates a feedback loop that enhances the initial model. Only M1 estimates the expected error of each location, though its explanatory power is limited (average $R^{2}=0.09$). All models predict the mean wealth of locations, but only M3 additionally predicts their standard deviation, and proposed two additional models to separate the effects of images and metadata.
2.3 Expected Trends
Before digging into the comparison of predictions between the three models of interest [2, 16, 8], it is crucial to acknowledge the discrepancies in the ground-truth data used to train them. Table 1 (top) outlines the specific years for which each approach utilized data for each country. This variation in time is significant because wealth and poverty are dynamic concepts that can fluctuate over time. To address this, rather than comparing directly the predicted values, we focus on the trends in predictions. We will achieve this by examining the trends of changes in Gross Domestic Product (GDP) per capita provided by the World Bank [27], as depicted in Figure 1 (top). By analyzing these trends, we can hypothesize that the comparison between RWI and IWI will mirror the trends observed in GDP. This approach allows us to account for potential discrepancies arising from the distinct time-frames of the datasets.
Additionally, we compare the changes in the heterogeneity of the predicted wealth distributions with each country’s official Gini index. Due to the lack of data from the World Bank [28], for the same years as the surveys in most countries, see Figure 1 (bottom), we present comparisons only for Uganda, where data is available, and South Africa, where all models used the same ground-truth.
Table 2: Prediction statistics. We show the number $N$ of target locations contained in the inferred poverty map, and the mean, standard deviation, and Gini coefficient of the predicted wealth scores for each model. RWI scores are linearly transformed to their IWI equivalents, see Equation 1.
Country M1 ($\hat{IWI}$) M2 (IWI) M3 (IWI)
$N$ Mean SD Gini $N$ Mean SD Gini $N$ Mean SD Gini
Sierra Leone 8435 29.12 4.92 8.79 13040 18.93 10.34 23.01 9881 21.87 6.09 11.54
Liberia 6189 28.62 3.71 6.72 16525 14.57 8.79 25.93 15597 23.35 3.74 7.75
Uganda 25404 41.35 3.06 3.92 27482 22.27 11.26 25.09 27791 41.83 5.21 5.95
Rwanda 3715 45.81 2.93 3.38 23323 24.40 14.51 28.96 1150 47.79 3.46 3.74
South Africa 39072 48.95 8.59 9.69 22687 61.14 18.10 16.85 3043 52.65 14.11 15.04
Gabon 1110 51.40 5.15 5.32 4835 21.89 8.86 17.93 1135 48.32 6.72 7.38
Table 3: Expected vs. obtained differences in wealth. The direction of change in wealth between M1 and M3 generally aligns with World Bank GDP trends, especially in SLE, LBR, UGA, RWA, and GAB. In contrast, the direction of change in wealth between M2 and M3 is only observed in RWA, with larger than expected differences. Most comparisons are significantly different ($p<0.001$, using a two-sample KS-test), except in SLE for M1 and M3. Highlighted values accurately capture the expected trend’s sign (+/-).
Country M1 ($\hat{IWI}$) vs. M3 ($IWI$) M2 ($IWI$) vs. M3 ($IWI$)
Exp.
IWI
Obt.
IWI
Exp.
Gini
Obt.
Gini
KS
Exp.
IWI
Obt.
IWI
Exp.
Gini
Obt.
Gini
KS
Sierra Leone 0.16 0.14 NA -2.75 0.74 0.01 -0.07 NA 11.47 0.62 ***
Liberia 0.04 0.10 NA -1.02 0.67 *** 0.04 -0.23 NA 18.19 0.83 ***
Uganda -0.04 -0.01 -1.7 -2.02 0.13 *** 0.0 -0.31 0.0 19.15 0.87 ***
Rwanda -0.05 -0.02 NA -0.36 0.34 *** -0.05 -0.32 NA 25.22 0.85 ***
South Africa 0.00 -0.04 0.0 -5.34 0.18 *** 0.0 0.07 0.0 1.81 0.30 ***
Gabon 0.11 0.03 NA -2.06 0.45 *** NA -0.38 NA 10.55 0.95 ***
Expected IWI Difference $=(GDP_{Y_{M*}}-GDP_{Y_{M3}})/(GDP_{Y_{M*}}+GDP_{Y_{M3}})$
Obtained IWI Difference $=(\mu_{IWI_{M*}}-\mu_{IWI_{M3}})/(\mu_{IWI_{M*}}+\mu_{IWI_{M3}})$
Expected Gini Difference $=(Gini_{Y_{M*}}-Gini_{Y_{M3}})$
Obtained Gini Difference $=(Gini_{M*}-Gini_{M3})$
3 Results
Here, we compare the inferred poverty maps produced by M1, M2, and M3. We first present the predicted overall wealth distributions. Due to the lack of ground truth in all places of these poverty maps, we compare the predicted distributions with official GDP and Gini coefficient trends to understand how well these predictions align with established economic indicators. Given that each study employs different spatial resolutions, our comparison then focuses on locations that appear in all three studies. Note that these analyses are based on the predicted poverty maps rather than test sets with ground truth. We trained all variants of M3 models and use the best model in each country based on their lowest RMSE scores (see details in Section 2.2) to infer the poverty maps. The inferences from the other two models are publicly available: M1 [18], M2 [15]. For a comparison of the reported predictive power in terms of R2 for these studies on their respective test sets, please refer to Appendix Table A1.
3.1 Predicted wealth distribution
Table 2 summarizes the poverty map inferences for each country and model. First, we observe that M2 generally identifies more locations (populated places) than M1 and M3. Second, after re-scaling RWI scores for a fair comparison (see Equation 1), we find that M2 predicts the lowest mean wealth in most cases, except in South Africa. In contrast, the mean wealth predicted by M1 and M3 between each country is very similar. Third, in terms of standard deviation, M1 predicts a more homogeneous wealth distribution, characterized by lower values of SD and Gini coefficient. M3 follows, and M2 exhibits the highest SD and Gini coefficients, indicating greater wealth inequality in its predictions.
(a) High-resolution
(b) Districts
(c) IWI distribution
Figure 2: IWI predictions in South Africa, ZAF (M3: CNNa+CB). All models were trained on the same ground-truth data (2016), thus, we expect similar results. Wealth predictions shown for M1, are linearly transformed into their IWI counterpart. Color bars display the average predicted IWI scores, centered at the mean of M3’s predictions for comparison. (a) Prediction at the original scale, 2.4 km grid-cells (M1), and OSM populated places (M2, M3). Both M2 and M3 include the Marion Island, omitted here for visualization purposes. (b) Aggregated mean wealth at the district level (administrative level 2). Compared to M3, M2 is consistently predicting higher wealth, while M1 predicts lower values. (c) Predicted wealth distribution (top: PDF, bottom: CDF). ZAF is the second country with the highest GDP in our study (see Figure 1), and this is reflected in the prediction by all models (e.g., $\mu>48$). Both M1 and M3 predict skewed distributions of wealth, while M2’s wealth follows an unusual multi-modal distribution. IWI poverty line ($35^{th}$ percentile, corresponding to poverty headcount ratios at \$1.25 a day [22]) is similar between M1 and M3, but differs for M2.
3.2 Expectation vs. Obtained Trends
We compared M1 and M2 to M3, focusing on their expected performance based on changes in GDP per captia per country over the years of each model’s data. [^1] Since M3 is used as a baseline for comparison with both M1 and M2, comparing M1 and M2 directly is unnecessary because their relative performance can be inferred from their respective comparisons to M3. Results are shown in Table 3. In Uganda, where M2 and M3 used the same ground-truth, we expect similar wealth predictions. However, M2 inferred lower and more heterogeneous wealth than M3. In South Africa, despite all models using the same ground-truth data, we found differing results: M1 predicted slightly lower wealth values than M3, and M2 predicted slightly higher values. These patterns are visualized in Figure 2 (Section 0.A.2 for other countries). In Liberia, models M1, M2, and M3 used ground-truth data from $2013$, $2016$, and $2019/2022$, respectively (Table 1). During this period, GDP per capita rose from $717.6$ in $2013$ to $722.1$ in $2016$, dropped to $665.9$ in $2019$, and increased to $754.5$ in $2022$ (Figure 1). The normalized GDP difference between $2013$ and $2019$ ($0.04$) suggests that M3’s IWI scores should be slightly lower than M1’s, and that is what we obtained ($0.1$). However, even though the same trend was expected between M2 and M3, M2’s IWI scores were unexpectedly lower ($-0.23$). In terms of Gini coefficients, M3 generally predicted more heterogeneous IWI scores than M1 and more homogeneous scores than M2. In Uganda, M2 indicated greater inequality than M3 ($19.15$), which in turn showed more inequality than M1 ($-2.02$). The same pattern was observed in South Africa, although similar poverty skewness was expected across models. Due to data limitations, we were unable to track Gini coefficient trends for the other countries.
(a) M1 vs. M2
(b) M1 vs. M3
(c) M2 vs. M3
Figure 3: IWI predictions in overlapping locations in South Africa, ZAF (M3: CNNa+CB). Left: Correlation between predictions of two different models. Right: Poverty map using only overlapping locations. Colors indicate differences in predicted values per model pair: (a) M1 vs. M2, (b) M1 vs. M3, and (c) M2 vs. M3. M2 generally overestimates wealth compared to M1 (mostly red in (a)) and M3 (mostly blue in (c)). Moreover, the wealth distribution in these locations is equally skewed. In (b), M1’s predictions are more homogeneous than M3’s, despite achieving relatively low RMSE.
Table 4: Prediction differences in overlapping locations. Due to the differing spatial resolutions between the models (grid-cells vs. OSM populated places), we compare predictions only for overlapping locations. We spatially joined locations from M1 and M3, and M2 and M3, defining overlap as places in study $i$ that are also in study $j$ within a distance of less than $500m$. As expected, M3 contains more overlapping locations $N_{o}$ with M2 than with M1, since both M2 and M3 use OSM populated places. Although M2’s predictions are more positively correlated with M3’s, M3’s predictions are closer to M1’s (low RMSE) than to M2’s (high RMSE), consistent with the overall distributions (Table 2). Highlighted values accurately capture the expected trend’s sign (+/-)
Country M1 ($\hat{IWI}$) vs. M3 ($IWI$) M2 ($IWI$) vs. M3 ($IWI$)
$N_{o}$
RMSE
Pearson
Corr.
IWI
Diff.
Gini
Diff.
$N_{o}$
RMSE
Pearson
Corr.
IWI
Diff.
Gini
Diff.
Sierra Leone 1125 9.72 0.69 *** -0.16 -1.08 7264 6.89 0.87 *** 0.08 10.34
Liberia 1332 6.59 0.53 *** -0.11 -0.51 14609 11.06 0.74 *** 0.26 15.66
Uganda 2799 3.69 0.73 *** -0.02 0.06 13057 19.58 0.82 *** 0.26 20.47
Rwanda 95 4.63 0.25 * -0.01 0.27 1603 21.77 0.43 *** 0.20 21.19
South Africa 314 12.98 0.59 *** -0.04 -6.02 2866 22.78 0.45 *** -0.13 0.52
Gabon 43 6.47 0.62 *** -0.03 -3.29 598 25.21 0.73 *** 0.32 18.06
IWI Difference $=(\mu_{IWI_{M*}}-\mu_{IWI_{M3}})/(\mu_{IWI_{M*}}+\mu_{IWI_{M3}})$
Gini Difference $=(Gini_{M*}-Gini_{M3})$
3.3 Overlapping locations
The previous results compared the overall predictions across all models, providing a general understanding of their performance relative to country-level GDP and Gini coefficient trends. Now, we focus on comparing the predictions for locations that overlap in all three studies. It is important to note that the three studies differ in the spatial resolution of their poverty maps: M1 uses grid cells, while M2 and M3 use OSM populated places (see details in Section 2.1). Therefore, we perform a spatial join to analyze only these overlapping locations, which we expect to perfectly reflect the anticipated trends.
As before, we use M3 as a baseline model to compare the differences across all models, with results presented in Table 4. Unexpectedly, the observed trends are not consistently maintained across most countries and models. Notable exceptions include Uganda and Rwanda, where M3 infers higher mean wealth than M1, and Sierra Leone and Liberia, where M3 infers lower mean wealth than M2, aligning with the expected GDP trends over time (see Table 3). In the specific case of South Africa, where all models were trained using the same ground truth data, M3’s mean wealth in these overlapping locations is higher than that inferred by the other models. Nevertheless, M2’s heterogeneity is closer to that M3, while M1 produced more homogeneous wealth scores, see details in Figure 3.
4 Conclusions
To assess the alignment between the IWI score predictions by three models—M1 (Chi et al., 2021 [2]), M2 (Lee and Braithwaite, 2022 [16]), and M3 (Espín-Noboa et al., 2023 [8])—we examined the differences in their underlying ground-truth data and compared predicted wealth trends against expected changes in GDP per capita. Overall, the differences between M1 and M3 generally met expectations. Sierra Leone, Liberia, and Gabon showed higher mean wealth in M1 compared to M3, while South Africa showed lower values, and Uganda and Rwanda had similar ones. Small differences suggested stable wealth in Liberia, Uganda, Rwanda, and South Africa over time (M1 $\rightarrow$ M3). However, M2’s predictions did not align as expected, indicating lower poverty levels in several countries, except in South Africa, where M2’s IWI scores were slightly higher than M3’s. In general, the wealth distribution predicted by M3 was slightly more heterogeneous than that inferred by M1, but it was more homogeneous compared to M2. This trend was particularly expected and observed in Uganda between M1 and M3, and in South Africa when comparing M2 and M3.
When comparing the trends in the locations that appear in all three studies (i.e., overlaps), the expected trends were not consistently observed in most countries. Nevertheless, predictions by M1 are the closest to those by M3 compared to M2, particularly in Liberia, Uganda, Rwanda, South Africa and Gabon. Similar to the overall distributions, M2’s predictions for this set of places produce the most heterogeneous wealth distribution (i.e., highest inequality). Given the lack of official Gini coefficients for the years of interest for most countries, we cannot evaluate whether these trends in heterogeneity are expected. However, in South Africa, we anticipated minimal differences as all models use the same ground truth data. Only M2 and M3 reflect this expected trend.
These findings indicate that the predictions by Chi et al. and Espín-Noboa et al. align with general GDP trends over time. Differences in their predictions are expected since each study used data from different periods. For instance, Sierra Leone’s GDP decline from 2013 to 2016 and 2019 (Figure 1) is reflected in the corresponding drop in (re-scaled) RWI (2013) and IWI (2016, 2019) scores (Table 2). Discrepancies with the overall predictions by Lee and Braithwaite, as well as among the overlapping locations, underscore the necessity for policymakers and African stakeholders to further audit these models and predicted maps to ensure their validity before deployment. While evaluations during training are important, they are insufficient to capture the overall validity of the inferred wealth indicators, as they subjectively focus on the model’s performance on the test data. Accurate poverty maps are essential for effective policy interventions and play a crucial role in addressing the United Nations’ first sustainable development goal: eradicating poverty world-wide.
{credits}
4.0.1 Acknowledgements
This work was funded by the SoBigData++ project (H2020-871042), Dataredux ANR project (ANR-19-CE46-0008), and the CHIST-ERA project SAI: FWFI 5205-N. The presented computational results have been achieved in part using the Vienna Scientifc Cluster (VSC). We thank the Data for Good team at Meta for insightful discussions during the first partner lightning talks, where early versions of this work were presented. We also appreciate the authors of the analyzed papers for their open data practices, which made this audit possible.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.
Appendix 0.A Appendix
In this appendix, we present the comparison of the three models, M1, M2, and M3, across the remaining five African countries considered in this study: Sierra Leone, Liberia, Uganda, Rwanda, and Gabon. This analysis supplements our main findings by providing additional insights into the performance and predictions of each model, highlighting the differences in wealth estimates.
Prediction data. Openly available here: M1 [18], M2 [15], and M3 [9].
0.A.1 Performance on test sets
Table A1: Reported validation R2 scores for each approach. M1 reports validation results on census data, as shown in Fig. S4 of the supplementary information [2]. M2 presents validation results for single-country training in Table 1 [16]. We retrain the M3 models for each country using a subset of the original features reported in [8], and report the performance (R2) of the best model, indicated by the lowest root-mean-square error, RMSE (see details in Section 2.2).
Country / R2 M1 M2 M3
Sierra Leone 83.0 89.82 80.7
Liberia - 89.3 76.3
Uganda - 89.02 78.5
Rwanda 82.0 87.41 67.5
South Africa - 73.03 49.5
Gabon - - 81.9
0.A.2 Overall predictions
(a) High-resolution
(b) Districts
(c) IWI distribution
Figure A1: IWI predictions in Sierra Leone (M3: CB). We expect M1 and M2 to produce wealthier maps than M3  (Table 3), with M1 being the wealthiest (Table 2). Wealth predictions shown for M1, are linearly rescaled. Color bars display the average predicted IWI scores, centered at the mean of M3’s predictions for comparison. (a) Prediction at the original scale. (b) Aggregated mean wealth at the administrative level 2. Compared to M3, M2 is consistently predicting lower wealth, while M1 predicts higher values (M1 and M3 fulfill the expectation). (c) Predicted wealth distribution (From top: PDF, CDF). IWI poverty line ($35^{th}$ percentile, corresponding to poverty headcount ratios at \$1.25 a day [22]) is similar between M2 and M3, but differs for M1.
(a) High-resolution
(b) Districts
(c) IWI distribution
Figure A2: IWI predictions in Liberia (M3: CB). We expect both M1 and M2 to produce slightly wealthier maps than M3  (Table 3). Wealth predictions shown for M1, are linearly rescaled. Color bars display the average predicted IWI scores, centered at the mean of M3’s predictions for comparison. (a) Prediction at the original scale. (b) Aggregated mean wealth at the administrative level 2. Compared to M3, M2 is consistently predicting lower wealth, while M1 predicts higher values (M1 and M3 fulfill the expectation). (c) Predicted wealth distribution (From top: PDF, CDF). IWI poverty line ($35^{th}$ percentile, corresponding to poverty headcount ratios at \$1.25 a day [22]) differs across models.
(a) High-resolution
(b) Districts
(c) IWI distribution
Figure A3: IWI predictions in Uganda (M3: CBw). We expect M3 to produce slightly wealthier maps than M1 and comparable to M2 (Table 3). Wealth predictions shown for M1, are linearly rescaled. Color bars display the average predicted IWI scores, centered at the mean of M3’s predictions for comparison. (a) Prediction at the original scale. (b) Aggregated mean wealth at the administrative level 2. Compared to M3, M2 is consistently predicting lower wealth, while M1 predicts slightly lower values (M1 and M3 fulfill the expectation). (c) Predicted wealth distribution (From top: PDF, CDF). IWI poverty line ($35^{th}$ percentile, corresponding to poverty headcount ratios at \$1.25 a day [22]) is similar between M1 and M3, but differs for M2.
(a) High-resolution
(b) Districts
(c) IWI distribution
Figure A4: IWI predictions in Rwanda (M3: CBw). We expect both M1 and M2 to produce slightly poorer poverty maps than M3 (Table 3). Wealth predictions shown for M1, are linearly rescaled. Color bars display the average predicted IWI scores, centered at the mean of M3’s predictions for comparison. (a) Prediction at the original scale. (b) Aggregated mean wealth at the administrative level 2. Compared to M3, both M2 and M1 predict lower wealth values. However, M2 is twice as poor than M1 (M1 and M3 fulfill the expectation). (c) Predicted wealth distribution (From top: PDF, CDF). IWI poverty line ($35^{th}$ percentile, corresponding to poverty headcount ratios at \$1.25 a day [22]) is similar between M1 and M3, but differs for M2.
(a) High-resolution
(b) Districts
(c) IWI distribution
Figure A5: IWI predictions in Gabon (M3: CNNa+CB). We expect M1’s poverty map to be wealthier than M3 (Table 3). Wealth predictions shown for M1, are linearly rescaled. Color bars display the average predicted IWI scores, centered at the mean of M3’s predictions for comparison. (a) Prediction at the original scale. (b) Aggregated mean wealth at the administrative level 2. Compared to M3, M1 predicts a wealther country (as expected). (c) Predicted wealth distribution (From top: PDF, CDF). IWI poverty line ($35^{th}$ percentile, corresponding to poverty headcount ratios at \$1.25 a day [22]) is similar between M1 and M3, but differs for M2.
0.A.3 Overlapping locations
(a) M1 vs. M2
(b) M1 vs. M3
(c) M2 vs. M3
Figure A6: IWI predictions in overlapping locations in Sierra Leone (M3: CB). Left: Correlation between the predictions of two models. Right: Overlapping locations. Colors indicate differences in predicted values per model pair. M1 generally overestimates wealth compared to M2 (blue in (a)) and M3 (blue in (b)), and predictions are more homogeneous. In (c), M2’s predictions are more heterogeneous than M3’s, despite achieving relatively low RMSE.
(a) M1 vs. M2
(b) M1 vs. M3
(c) M2 vs. M3
Figure A7: IWI predictions in overlapping locations in Liberia (M3: CB). Left: Correlation between the predictions of two models. Right: Overlapping locations. Colors indicate differences in predicted values per model pair. M1 generally overestimates wealth compared to M2 (blue in (a)) and M3 (blue in (b)), and predictions are more homogeneous. In (c), M2’s predictions are more heterogeneous than M3’s.
(a) M1 vs. M2
(b) M1 vs. M3
(c) M2 vs. M3
Figure A8: IWI predictions in overlapping locations in Uganda (M3: CBw). Left: Correlation between the predictions of two models. Right: Overlapping locations. Colors indicate differences in predicted values per model pair. M1 generally overestimates wealth compared to M2 (blue in (a)) and M3 (blue in (b)). Predictions from M1 are more homogeneous than M2 and equally skewed as M3. In (c), M2’s predictions are more heterogeneous than M3’s.
(a) M1 vs. M2
(b) M1 vs. M3
(c) M2 vs. M3
Figure A9: IWI predictions in overlapping locations in Rwanda (M3: CBw). Left: Correlation between the predictions of two models. Right: Overlapping locations. Colors indicate differences in predicted values per model pair. M2 generally underestimates wealth compared to M1 (blue in (a)) and M3 (red in (c)). Predictions from M1 are more homogeneous than M2 and equally skewed as M3. In (c), M2’s predictions are more heterogeneous than M3’s.
(a) M1 vs. M2
(b) M1 vs. M3
(c) M2 vs. M3
Figure A10: IWI predictions in overlapping locations in Gabon (M3: CNNa+CB). Left: Correlation between the predictions of two models. Right: Overlapping locations. Colors indicate differences in predicted values per model pair. M2 generally underestimates wealth compared to M1 (blue in (a)) and M3 (red in (c)). Predictions from M1 are more homogeneous than M2 and M3. In (c), M2’s predictions are more heterogeneous than M3’s.